# On biquandle-based invariant of immersed surface-links and on Yoshikawa oriented fifth move

### v1

@@ Abstract.
We resolve an open problem posed by Yoshikawa, showing that the fifth oriented move in his list cannot be reproduced by any finite sequence of the other nine moves and planar isotopies. Our proof introduces a link-type semi-invariant that remains unchanged under all moves except the fifth; contrasting values on two equivalent diagrams force the move’s independence. Second, we extend the algebraic toolkit for immersed surface-links. After revisiting the banded-unlink description of immersed surfaces and the twelve local moves that relate their diagrams, we develop a colouring theory based on biquandles. By assigning elements of a biquandle to diagram arcs according to local rules, we obtain a counting invariant of immersed surfaces up to isotopy.
@@ 1. Introduction
The paper tackles two central problems in the diagrammatic study of oriented surfaces immersed in $\mathbb{R}^{4}$
First, it settles an open question about Yoshikawa moves, the ten local moves that generate all diagrammatic isotopies of oriented surface-links. Previous work had proved the independence of every move except the oriented fifth move. We construct a semi-invariant derived from admissible marked-graph diagrams whose value is unchanged by every move in the generating set of other moves. Exhibiting two diagrams of the same surface-link with different values.
Second, the paper enriches the algebraic toolkit for immersed (i.e. generically self-intersecting) surface-links. We revisit the singular marked-graph (or “banded unlink”) presentation, extending the catalogue of oriented planar moves so that any two diagrams of equivalent immersed surfaces differ by the twelve types of moves. We equip an immersed surface-link with a biquandle structure. Each semi-arc of a singular marked diagram is coloured by elements of a biquandle according to local rules that respect crossings, markers and singular points. Because the biquandle axioms are tuned to the $\Gamma$-moves, the number of colourings is an invariant of the underlying surface. The authors compute this new biquandle colouring invariant for an illustrative example, obtaining two admissible colourings with a four-element biquandle.
Taken together, the independence proof sharpens our understanding of the move calculus, while the colouring invariant supplies fresh algebraic data for distinguishing immersed surfaces—advancing both the combinatorial and algebraic sides of four-dimensional knot theory.
@@ 2. Independence of the Yoshikawa oriented fifth move
It is known that the set of ten types of moves $\mathcal{G}=\{\Gamma_{1},\ldots,\Gamma_{8},\Gamma_{4}^{\prime},\Gamma_{6}^{\prime}\}$ (presented in Fig. 1 and  2), called oriented Yoshikawa moves, is a generating set of moves that relates two marked graph diagrams (modulo a planar isotopy) presenting equivalent oriented surface-links. In ([JKL13], [JKL15]) it is shown that any Yoshikawa move from the set $\mathcal{G}\backslash\{\Gamma_{5},\Gamma_{8}\}$ is independent from the other nine types. It is an open problem (see [JKL15], [Oht16]) whether the move $\Gamma_{5}$ is independent of the other Yoshikawa moves from the set $\mathcal{G}$.
@@ 3. Singular banded unlinks
By general position theorems for maps any smooth map $f:X^{n}\to Y^{2n}$ can be approximated homotopically with arbitrary accuracy by an immersion described above.
@@ 4. Biquandle structure
A generalization of quandles (called biquandles) was introduced in [KauRad03]. A biquandle is an algebraic structure with two binary operations satisfying certain conditions which can be presented by semi-arcs of links (or semi-sheets of surface-links) as its generators modulo oriented Reidemeister moves (or Roseman moves). In [CES04], J. S. Carter, M. Elhamdadi and M. Saito introduced and used cocycles to define invariants via colorings of link diagrams by biquandles and a state-sum formulation.
@@
We can enumerate that #$\#Col_{X_{T}}(D_{T})=3$, the one of the admissible coloring is shown in Figure 13, the other admissible coloring is obtained from the previous by exchanging colors $1$ with $3$, and the third one is monochromatic by element $2$.
@@
[]michal_031(5cm)
Figure 13. An admissible coloring of the diagram $D_{T}$.

### v2

@@ Abstract.
We resolve an open problem showing that the Yoshikawa’s fifth oriented move in his list cannot be reproduced by any finite sequence of the other nine moves and planar isotopies. Our proof introduces a link-type semi-invariant that remains unchanged under all moves except the fifth; contrasting values on two equivalent diagrams force the move’s independence. Second, we extend the algebraic toolkit for immersed surface-links. After revisiting the banded-unlink description of immersed surfaces and the twelve local moves that relate their diagrams, we develop a coloring theory based on biquandles. By assigning elements of a biquandle to diagram arcs according to local rules, we obtain a counting invariant of immersed surfaces up to isotopy.
@@ 1. Introduction
The paper tackles two central problems in the diagrammatic study of oriented surfaces immersed in $\mathbb{R}^{4}$.
First, it settles an open question about Yoshikawa’s moves, the ten local moves that generate all diagrammatic isotopies of oriented surface-links. Previous work had proved the independence of every move except the oriented fifth move. We construct a semi-invariant derived from admissible marked-graph diagrams whose value is unchanged by every move in the generating set of other moves. Exhibiting two diagrams of the same surface-link with different values.
Second, the paper enriches the algebraic toolkit for immersed (i.e. generically self-intersecting) surface-links. We revisit the singular marked-graph (or “banded unlink”) presentation, extending the catalog of oriented planar moves so that any two diagrams of equivalent immersed surfaces differ by the twelve types of moves. We equip an immersed surface-link with a biquandle structure. Each semi-arc of a singular marked diagram is colored by elements of a biquandle according to local rules that respect crossings, markers and singular points. Because the biquandle axioms are tuned to the $\Gamma$-moves, the number of colorings is an invariant of the underlying surface. The authors compute this new biquandle coloring invariant for an illustrative example, obtaining two admissible colorings with a four-element biquandle.
Taken together, the independence proof sharpens our understanding of the move calculus, while the coloring invariant supplies fresh algebraic data for distinguishing immersed surfaces, advancing both the combinatorial and algebraic sides of four-dimensional knot theory.
@@
Acknowledgements
”This research was funded in whole or in part by NCN 2023/07/X/ST1/00157. For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.”
@@ 2. Independence of the Yoshikawa oriented fifth move
It is known that the set of ten types of moves $\mathcal{G}=\{\Gamma_{1},\ldots,\Gamma_{8},\Gamma_{4}^{\prime},\Gamma_{6}^{\prime}\}$ (presented in Fig. 1 and  2), called oriented Yoshikawa moves ([Yos94]), is a generating set of moves that relates two marked graph diagrams (modulo a planar isotopy) presenting equivalent oriented surface-links. In ([JKL13], [JKL15]) it is shown that any Yoshikawa move from the set $\mathcal{G}\backslash\{\Gamma_{5},\Gamma_{8}\}$ is independent of the other nine types. It is an open problem (see [JKL15], [Oht16]) whether the move $\Gamma_{5}$ is independent of the other Yoshikawa moves from the set $\mathcal{G}$.
@@ 3. Singular banded unlinks
By general position theorems for maps, any smooth map $f:X^{n}\to Y^{2n}$ can be approximated homotopically with arbitrary accuracy by an immersion described above.
@@ 4. Biquandle structure
A generalization of quandles (called biquandles) was introduced in [KauRad03]. A biquandle is an algebraic structure with two binary operations satisfying certain conditions which can be presented by semi-arcs of links (or semi-sheets of surface-links) as its generators modulo oriented Reidemeister moves (or Roseman moves). In [CES04], J. S. Carter, M. Elhamdadi, and M. Saito introduced and used cocycles to define invariants via colorings of link diagrams by biquandles and a state-sum formulation.
@@
We can enumerate that #$\#Col_{X_{T}}(D_{T})=5$, two of the admissible coloring is shown in Figure 13, the other two admissible colorings are obtained from the previous by exchanging colors $1$ with $3$, and the fifth one is monochromatic, colored by element $2$.
@@
[]michal_032(9cm)
Figure 13. Admissible colorings of the diagram $D_{T}$.